# Bayesian spatio-temporal weighted regression for integrating missing and misaligned environmental data

### v1

@@ 4.1 Data
The dataset comprises fine particulate matter (PM2.5) measurements from 67 air quality monitoring stations across Northern California between 1st June 2020 and 31st October 2020, coinciding with the region’s peak wildfire season. Ground-based data were augmented with satellite data of aerosol optical depth (AOD) from the Multi-Angle Implementation of Atmospheric Correction (MAIAC) algorithm, which provides high-resolution (1̃ km x 1 km) gridded observations from the Moderate Resolution Imaging Spectroradiometer (MODIS) instruments aboard NASA’s Terra and Aqua satellites [Lyapustin et al., 2018]. To account for wildfire smoke influence on PM2.5 concentrations, smoke plume data were obtained from the National Oceanic and Atmospheric Administration’s (NOAA) Hazard Mapping System (HMS), which delineates smoke plume extent based on satellite imagery [NOAA Satellite and Information Service, 2025]. Together these data enabled a comprehensive characterization of spatiotemporal variation in PM2.5 and its association with wildfire smoke during the 2020 fire season in Northern California.
Figure 3: Map of Northern California study region on two consecutive days in June 2020 showing: (a) PM2.5 monitoring locations (blue dots) with 20 km buffers around them (shaded blue circles) with 1 km x 1 km MAIAC aerosol optical depth (AOD) grid (pink to red); (b) zoom in of one PM2.5 monitor on a day with nearly completely observed AOD within 20 km; and (c) zoom in of one PM2.5 monitor on the next day with significant missing AOD within 20 km
@@ 4.2 Model Specifications and Estimation
Building on the baseline model, the functional mean model augments the intercept by a temporally varying mean function $\beta_{0}(t)$ represented with B-spline basis functions $f_{k}(t)$ [Wood, 2017]. We used $K=10$ equally spaced knots to capture bi-weekly trends, and retained the same priors as in the baseline model.
In the remaining models we included a varying coefficient for wildfire smoke. Let $I(\mathbf{s},t)$ indicate a smoke plume event (1 = satellite-derived wildfire smoke plume intersected a PM2.5 monitor $Y(\mathbf{s},t)$ at location $s$ on day $t$, 0 = otherwise). We incorporated this indicator to the models: (1) at the mean level where $\beta_{0}(\mathbf{s},t)=\sum_{k}b_{jk}f_{k}(t)+I(\mathbf{s},t)$, (2) as a varying coefficient where the effect of the weighted predictor of AOD varied by plume status $\beta_{1}(\mathbf{s},t)=\beta_{10}+\beta_{11}I(\mathbf{s},t)$, and (3) by fixing $\xi_{2}$ at the posterior median from the functional mean model and allowing the spatial scale to depend on smoke plume status. This was motivated by [Zhang, 2004], who showed that range and scale parameters in covariance functions are not consistently estimable and require stabilization.
The full model includes all features above and allows the spatial extent over which AOD influences PM2.5 to vary based on smoke plumes via $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}\,I(\mathbf{s},t)+\nu,\qquad\nu\sim N(0,\sigma_{\nu}^{2})$. To help estimation, when $I(\mathbf{s},t)=0$, we placed a strong point-mass prior on $\xi_{1}(\mathbf{s},t)$ centered at 2.0 for $r=10$ and 1.0 for $r=20$ (values obtained under the varying coefficient model). Under this approach, we effectively only estimated the effect of $I(\mathbf{s},t)$ through $\alpha_{1}$. The MCMC algorithms used for inference can be found under Appendix D. Convergence was assessed using the Gelman-Rubin test and effective sample size. Each model was run for 25000 MCMC iterations, with the first 12500 samples used as warm-up and retaining the latter 12500 of the samples for inference. As a benchmark, we also fit a collocated linear model where we aggregated and averaged AOD data within a 5 km radius around each station on each day.
@@
Full $\sum_{k}b_{0k}\,f_{k}(t)$ $\gamma\,I(\mathbf{s},t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\log\!\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}$ $\alpha_{1}I(\mathbf{s},t)+\nu,$ $\nu\sim N(0,\sigma_{\nu}^{2})$ $\xi_{2}$ fixed at posterior median
@@
 Following [Zhang, 2004], $\xi_{2}$ is fixed at the posterior median from the functional mean model to stabilize range estimation. In all models, $\beta_{1}(\mathbf{s},t)$ multiplies $W(\mathbf{s},t)$, the kernel-weighted AOD predictor defined in Eq.(1), which averages $X(\mathbf{u},t-l)$ over $\mathcal{B}_{r}(\mathbf{s})\times[0,q]$ using normalized exponential kernels in space and time. The collocated baseline model uses the AOD average $\overline{X}_{5\text{km}}(\mathbf{s},t)$ in 5 km buffer instead.
@@ 4.3 Results
The performance of each method is assessed using the root-mean square error (RMSE) metric, percentage of variance explained by the model (R2) and coverage of the 95% prediction interval for stations that were kept out-of-sample.
The baseline collocated linear model produced an RMSE of 25.8, $R^{2}=0.184$ and coverage 0.941. To fit this model, 407 observations were removed due to unpaired PM2.5-AOD observations.
Out-of-sample predictive performance of the competing models is reported in Table 2. Overall, the smaller buffer radius generally improved predictive accuracy, with the varying coefficient model benefitting less from this adjustment. In particular, the full model performs noticeably worse when the radius is set to $r=20$, suggesting that the added complexity at this radius may induce overfitting and reduce generalizability. However, we observed that the error is largely driven by overinflated predictions at a single location; when that location is removed, the model performance improves to RMSE = 19.3, R2 = 0.455, and coverage = 0.942.
By contrast, when the radius was restricted to $r=10$, the full model achieved the best balance across evaluation metrics, outperforming the alternative specifications. These findings highlight the sensitivity of model performance to the choice of buffer radius and emphasize the importance of carefully selecting this parameter in practice.
@@
Model RMSE $\boldsymbol{R^{2}}$ Coverage RMSE $\boldsymbol{R^{2}}$ Coverage
Baseline 32.5 0.230 0.956 35.1 0.227 0.957
Functional mean 19.6 0.464 0.940 21.0 0.410 0.942
Varying coefficient 21.0 0.362 0.944 19.9 0.400 0.944
Full 19.0 0.454 0.940 63.3 0.037 0.940
Collocated baseline model: RMSE = 25.8, $R^{2}=0.184$, Coverage = 0.941.
To gain further insight into the spatio-temporal dependencies between PM2.5 and AOD, Figure 4 illustrates the estimated surface $\hat{\omega}(h,l)=\hat{\beta}_{1}\,\kappa_{1}(h;\hat{\xi}_{1}(\mathbf{s},t))\,\kappa_{2}(l;\hat{\xi}_{2}(\mathbf{s},t)),$ for the two best-performing models: the functional mean ST-Exp model (Figure 4(a)) and the full ST-Exp model (Figures 4(b) and 4(c)), both with $r=10$. We used the posterior medians of the MCMC samples to construct the surface $\widehat{\omega}(h,l)$. Figure 4(b) shows the estimated effect of AOD on PM2.5 when no smoke plume was present ($I(\mathbf{s},t)=0$), whereas Figure 4(c) shows the effect under plume conditions ($I(\mathbf{s},t)=1$). The results indicate that incorporating wildfire smoke information alters the spatio-temporal weighting: in the functional mean model, the AOD effect was distributed more evenly across the buffer region, while in the full model, both for $I(\mathbf{s},t)=0$ and $I(\mathbf{s},t)=1$, the effect decayed more rapidly over space and time but attained larger magnitudes in the immediate neighbourhood of an observation. Taken together, these results suggest that fire events amplify the local influence of AOD on PM2.5, underscoring the importance of accounting for such events when constructing accurate exposure maps.
@@
(b) Full model in the absence of fire plumes
(c) Full model in the presence of fire plumes
Figure 4: Estimates of the surface explaining the spatio-temporal dependencies between PM2.5 and AOD at varying spatial and time lags ($r=10$).
Figure 5: Predictions of PM2.5 at out-of-sample locations over time. The top panel compares the collocated linear model with the functional mean model ($r=10$). The bottom panel compares the functional mean model with the full model ($r=10$). As model complexity increases, uncertainty quantification of the predictions improves with narrower prediction intervals capturing peaks.
Figure 5 compares out-of-sample temporal predictions and 95% posterior prediction intervals at held-out monitoring stations. The collocated linear benchmark model attained reasonable point predictions but produced very wide intervals, especially during sharp concentration peaks. By contrast, the functional mean model has substantially narrower interval widths while maintaining coverage, providing more precise uncertainty quantification for PM2.5 concentrations across both peak episodes and stable periods. Figure 5(b) further shows that incorporating wildfire information improves temporal calibration. The full model produced narrower prediction intervals that still captured the observed peaks, highlighting the importance for wildfire influences in PM2.5 estimation.
Figure 6 presents 5 km resolution gridded maps of estimated PM2.5 derived from 1 km resolution AOD for three days in September during peak wildfire season, using the functional mean model. The approach effectively captured both spatial gradients and day-to-day temporal patterns, producing smooth, coherent variation across space and time, particularly where the August Complex wildfire occurred.
Figure 6: Spatial mapping of PM2.5 using the functional mean model for three consecutive days in September 2020.
@@ 5 Discussion
Handling spatial misalignment traditionally relies on change-of-support or hierarchical modeling to reconcile point- and areal-level data [Gotway and Young, 2002]. In point-to-point and point-to-grid misalignment, common practice is to average data from the misaligned source over buffer regions around a target point to align (i.e. forced collocation). This approach can clearly induce misclassification for several reasons including choice of the buffer radius. Furthermore, misalignment may be present due to missing data. Our contribution is to embed the missing data problem inside a normalized spatio-temporal weighting operator whose kernel parameters can vary with covariates, coupled with a numerically stable quadrature that respects missingness and irregular supports. This yielded a unified route to estimation and uncertainty quantification without a separate imputation stage, while remaining computationally tractable.
@@
The weighted kernel-averaged predictor regression model we developed required careful specification of the spatial and temporal lag windows $(r,q)$ for valid inference and accurate prediction. Our theoretical results quantify how misspecification of $(r,q)$ propagated to bias in the mean, which in turn lead to biased parameter estimates and poorer predictive performance. Our proposed quadrature scheme that uses Voronoi-based spatial weights and irregular time steps to estimate the spatio-temporal integral, enabled estimation under extensive missing data and spatio-temporal misalignment. Unlike common strategies that first gap-fill (interpolate) and then regress on the complete interpolated surface (e.g. [Li and Heap, 2014]), our approach builds the weighting and uncertainty handling directly into the likelihood via $W(\mathbf{s},t)$, allowing propagation of uncertainty from the available data to the predictions.
@@
Missingness and and spatial misalignment are common issues in many environmental modeling applications. In the satellite AOD–PM2.5 literature, missing AOD due to clouds, snow, high aerosol loadings, or retrieval failures is widespread, and many studies use explicit gap filling via interpolation or ensembles before regression [Xiao et al., 2021, Sorek-Hamer et al., 2020, Li et al., 2020]. When we applied our approach to estimate PM2.5 from satellite AOD over Northern California in 2020, which had an intense wildfire season, the weighted predictor model captured both spatial and temporal variation and yielded narrower, better-calibrated prediction intervals than a collocated linear baseline, with more than a 50% increase in R2. Incorporating smoke plume information altered the estimated weighting surface and improved temporal calibration around concentration peaks. The proposed estimator handled missing and misaligned data effectively because it aggregated over whatever nearby $(\mathbf{u},l)$ were observed, rather than requiring a fully gridded covariate. When no informative neighbours existed within $(r,q)$, the observation is flagged rather than imputed, avoiding potentially biased reconstructions.
@@ Appendix A Voronoi Tessellation
Appendix B Simulation Study
Figure 8: Examples of simulated $X(s,t)$ fields.
@@ Appendix C Proof of Proposition 1
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{0}^{q}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell.$
@@
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta\int_{\mathcal{B}_{r}(\mathbf{s})}\int_{\min(q,\tilde{q})}^{\max(q,\tilde{q})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\ell\,d\mathbf{u}.$
@@
    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{2}$ minus the integral over $A_{1}$ in magnitude, and is therefore bounded above by the integral of the integrand over

### v2

@@ 4.1 Data
The dataset comprises of fine particulate matter (PM2.5) measurements from 67 air quality monitoring stations across Northern California between 1st June 2020 and 31st October 2020, coinciding with the region’s peak wildfire season. Ground-based data were linked with satellite data of aerosol optical depth (AOD) from the Multi-Angle Implementation of Atmospheric Correction (MAIAC) algorithm, which provides high-resolution (1̃ km x 1 km) gridded observations from an algorithm applied to observations from the Moderate Resolution Imaging Spectroradiometer (MODIS) instrument aboard NASA’s Terra and Aqua satellites (lyapustin2018). To account for wildfire smoke influence on PM2.5 concentrations, we included smoke plume data from the National Oceanic and Atmospheric Administration’s (NOAA) Hazard Mapping System (HMS), which delineates smoke plume extent based on satellite imagery (NOAA_HMS_Smoke). Together these data enabled the characterization of spatiotemporal variation in PM2.5 and its association with wildfire smoke during the 2020 fire season in Northern California.
Figure 3: Map of Northern California study region on two consecutive days in June 2020 showing: (a) PM2.5 monitoring locations (blue dots) with 20 km buffers around them (shaded blue circles) with 1 km x 1 km MAIAC aerosol optical depth (AOD) grid (pink to red); (b) zoom in of one PM2.5 monitor on a day with nearly completely observed AOD within 20 km; and (c) zoom in of one PM2.5 monitor on the next day with significant missing AOD within 20 km.
@@ 4.2 Model Specifications and Estimation
Building on the baseline model, the functional mean model augments the intercept by a temporally varying mean function $\beta_{0}(t)$ represented with B-spline basis functions $f_{k}(t)$ (wood_generalized_2017). We used $K=10$ equally spaced knots to capture bi-weekly trends, and retained the same priors as in the baseline model.
In the remaining models we included a varying coefficient for wildfire smoke. Let $I(\mathbf{s},t)$ indicate a smoke plume event (1 = satellite-derived wildfire smoke plume intersected a PM2.5 monitor $Y(\mathbf{s},t)$ at location $s$ on day $t$, 0 = otherwise). We incorporated this indicator to the models: (1) at the mean level where $\beta_{0}(\mathbf{s},t)=\sum_{k}b_{0k}f_{k}(t)+\gamma I(\mathbf{s},t)$ assuming a normal prior for $\gamma$, (2) as a varying coefficient where the effect of the weighted predictor of AOD varied by plume status $\beta_{1}(\mathbf{s},t)=\beta_{10}+\beta_{11}I(\mathbf{s},t)$. We fix $\xi_{2}$ at the posterior median from the functional mean model and allowing the effect of the kernel-averaged AOD to depend on smoke plume status. This was motivated by zhang_inconsistent_2004, who showed that range and scale parameters in covariance functions are not consistently estimable and require stabilization.
The full model includes the varying coefficient specification above and allows the spatial extent over which AOD influences PM2.5 to vary based on smoke plumes via $\log\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}+\alpha_{1}\,I(\mathbf{s},t)+\nu,\qquad\nu\sim N(0,\sigma_{\nu}^{2})$. To help estimation, when $I(\mathbf{s},t)=0$, we placed a strong point-mass prior on $\xi_{1}(\mathbf{s},t)$ centered at the posterior median of the samples obtained under the varying coefficient model. We assume normal priors for $\mathbf{\alpha}$ and inverse-gamma for $\sigma_{\nu}^{2}$. Under this approach, we effectively only estimated the effect of $I(\mathbf{s},t)$ through $\alpha_{1}$. The MCMC algorithms used for inference can be found under Appendix D in the Supplementary Material. Convergence was assessed using the Gelman-Rubin test. Each model was run for 25000-50000 MCMC iterations, with the first 50% of the samples used as warm-up and retaining the latter half for inference. As a benchmark, we also fit a collocated linear model where we aggregated and averaged AOD data within a 5 km radius around each station on each day.
@@
Full $\sum_{k}b_{0k}\,f_{k}(t)$ $\beta_{1}(\mathbf{s},t)=\beta_{10}$ $\beta_{11}I(\mathbf{s},t)$ $\log\!\{\xi_{1}(\mathbf{s},t)\}=\alpha_{0}$ $\alpha_{1}I(\mathbf{s},t)+\nu,$ $\nu\sim N(0,\sigma_{\nu}^{2})$ $\xi_{2}$ fixed at posterior median
@@
 Following zhang_inconsistent_2004, $\xi_{2}$ is fixed at the posterior median from the functional mean model to stabilize range estimation. In all models, $\beta_{1}(\mathbf{s},t)$ multiplies $W(\mathbf{s},t)$, the kernel-weighted AOD predictor defined in Eq.(1), which averages $X(\mathbf{u},t-l)$ over $\mathcal{B}_{r}(\mathbf{s})\times[0,q]$ using normalized exponential kernels in space and time. The collocated linear benchmark model uses the AOD average $\overline{X}_{5\text{km}}(\mathbf{s},t)$ in 5 km buffer instead.
@@ 4.3 Results
The performance of each method is assessed using the root-mean square error (RMSE) metric, continuous ranked probability score (CRPS) to evaluate the accuracy of predictive forecasts (gneiting2007strictly) and the coverage of the 95% prediction interval for stations that were kept out-of-sample.
The collocated linear model produced an RMSE of 25.8, CRPS 6.89 and coverage 0.941. To fit this model, 407 observations were removed due to unpaired PM2.5-AOD observations.
Out-of-sample predictive performance of the competing models is reported in Table 2. Overall, the larger radius around each site (r = 20 km) had better predictive accuracy (lower RMSE and CRPS), except for the full model, which performed better with r = 10 km. When $r=20$ the varying coefficient model achieved the best balance across evaluation metrics, outperforming the alternative model specifications.
@@
Model RMSE CRPS Coverage RMSE CRPS Coverage
Baseline 36.3 8.49 0.957 36.1 8.22 0.953
Functional mean 19.6 6.32 0.942 20.1 6.21 0.945
Varying coefficient 23.5 6.97 0.942 19.1 6.07 0.946
Full 19.6 6.22 0.943 20.1 6.34 0.944
[…]
To gain further insight into the spatio-temporal dependencies between PM2.5 and AOD, Figure 4 illustrates the estimated surface $\hat{\omega}(h,l)=\hat{\beta}_{1}\,\kappa_{1}(h;\hat{\xi}_{1}(\mathbf{s},t))\,\kappa_{2}(l;\hat{\xi}_{2}(\mathbf{s},t)),$ for the three best-performing models: the functional mean model (Figure 4(a)), the varying coefficient model (Figures 4(b) and 4(c)), both with $r=20$, and the full model (Figures 4(d) and 4(e)) with $r=10$. We used the posterior medians of the MCMC samples to construct the surface $\widehat{\omega}(h,l)$. Figures 4(b) and  4(d) show the estimated effect of AOD on PM2.5 when no smoke plume was present ($I(\mathbf{s},t)=0$), whereas Figure 4(c) and  4(e) show the effect under plume conditions ($I(\mathbf{s},t)=1$). The results indicate that incorporating wildfire smoke information alters the spatio-temporal weighting. The varying coefficient model is able to isolate the effect of AOD under different plume conditions. In the presence of fire plumes, the model estimates higher levels of PM2.5 with $\hat{\gamma}\in[0.461,0.587]$ on average while the effect of the kernel-averaged AOD being smaller and more localised in its immediate neighbourhood. The full model captures this differently, the magnitude of the effect of the weighted AOD on PM2.5 is higher in the presence of smoke plumes. These results suggest that fire events amplify the local influence of AOD on PM2.5, underscoring the importance of accounting for such events when constructing accurate exposure maps.
@@
($r=20$)
(b) Varying coefficient model in the
absence of fire plumes ($r=20$)
(c) Varying coefficient model in the presence of fire plumes ($r=20$)
(d) Full model in the absence of
fire plumes ($r=10$)
(e) Full model in the presence of fire plumes ($r=10$)
Figure 4: Estimates of the surface explaining the spatio-temporal dependencies between PM2.5 and AOD at varying spatial and time lags.
Figure 5: Predictions of PM2.5 at out-of-sample locations over time. The top panel compares the collocated linear model with the baseline model ($r=20$). The bottom panel compares the baseline model with the varying coefficient model ($r=20$). As model complexity increases, uncertainty quantification of the predictions improves with narrower prediction intervals capturing peaks.
Figure 5 compares out-of-sample temporal predictions and 95% posterior prediction intervals at held-out monitoring stations. The collocated linear benchmark model attained reasonable point predictions but produced very wide intervals, especially during sharp concentration peaks. By contrast, the baseline model has substantially narrower interval widths while maintaining coverage, providing more precise uncertainty quantification for PM2.5 concentrations across both peak episodes and stable periods. Figure 5(b) further shows that incorporating wildfire information improves temporal calibration. The full model produced narrower prediction intervals that still captured the observed peaks, highlighting the importance for wildfire influences in PM2.5 estimation.
Figure 6 presents 1 km resolution gridded maps of estimated PM2.5 derived from 1 km resolution AOD for three days in September during peak wildfire season, using the functional mean model. The approach effectively captured both spatial gradients and day-to-day temporal patterns, producing smooth, coherent variation across space and time, particularly where the August Complex wildfire occurred. An advantage of our approach is it enables fine-resolution predictions even when AOD is missing.
Figure 6: Spatial mapping of PM2.5 using the full model ($r=10$) for three consecutive days in September 2020.
@@ 5 Discussion
Handling spatial misalignment traditionally relies on change-of-support or hierarchical modeling to reconcile point- and areal-level data (gotway_young_2002). In point-to-point and point-to-grid misalignment, common practice is to average data from the misaligned source over buffer regions around a target point to align (i.e. forced collocation). This approach can clearly induce misclassification for several reasons including choice of the buffer radius. Furthermore, misalignment may be present due to missing data. Our contribution is to embed the missing data problem inside a normalized spatio-temporal weighting operator whose kernel parameters can vary with covariates, coupled with a numerically stable quadrature that respects missingness and irregular supports. This yielded a unified route to estimation and uncertainty quantification without a separate imputation stage, while remaining computationally tractable.
@@
The weighted kernel-averaged predictor regression model we developed required careful specification of the spatial and temporal lag windows $(r,q)$ for valid inference and accurate prediction. Our theoretical results quantify how misspecification of $(r,q)$ propagated to bias in the mean, which in turn lead to biased parameter estimates and poorer predictive performance. Our proposed quadrature scheme that uses Voronoi-based spatial weights and irregular time steps to estimate the spatio-temporal integral, enabled estimation under extensive missing data and spatio-temporal misalignment. Unlike common strategies that first gap-fill (interpolate) and then regress on the complete interpolated surface (e.g. li2014), our approach builds the weighting and uncertainty handling directly into the likelihood via $W(\mathbf{s},t)$, allowing propagation of uncertainty from the available data to the predictions.
@@
Missingness and and spatial misalignment are common issues in many environmental modeling applications. In the satellite AOD–PM2.5 literature, missing AOD due to clouds, snow, high aerosol loadings, or retrieval failures is widespread, and many studies use explicit gap filling via interpolation or ensembles before regression (xiao2021; sorek-hamer2020; li2020). When we applied our approach to estimate PM2.5 from satellite AOD over Northern California in 2020, which had an intense wildfire season, the weighted predictor model captured both spatial and temporal variation and yielded narrower, better-calibrated prediction intervals than a collocated linear baseline. Incorporating smoke plume information altered the estimated weighting surface and improved temporal calibration around concentration peaks. The proposed estimator handled missing and misaligned data effectively because it aggregated over whatever nearby $(\mathbf{u},l)$ were observed, rather than requiring a fully gridded covariate. When no informative neighbours existed within $(r,q)$, the observation is flagged rather than imputed, avoiding potentially biased reconstructions.
@@ Appendix A Voronoi Tessellation
Figure  7 shows the Voronoi tessellation over a buffer of 20km for different days. On days where there is missing AOD data, the region is weighted more heavily due to a larger Voronoi cell.
Appendix B Simulation Study
@@ Appendix C Proof of Proposition 1
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta_{1}\int_{0}^{q}\int_{\mathcal{A}_{r,\tilde{r}}(\mathbf{s})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\mathbf{u}\,d\ell.$
@@
    $\bigl|\mu(\mathbf{s},t)-\tilde{\mu}(\mathbf{s},t)\bigr|\leq\beta_{1}\int_{\mathcal{B}_{r}(\mathbf{s})}\int_{\min(q,\tilde{q})}^{\max(q,\tilde{q})}\kappa_{1}(\mathbf{s}-\mathbf{u})\,\kappa_{2}(t-\ell)\,X(\mathbf{u},t-\ell)\,d\ell\,d\mathbf{u}.$
@@
    Thus the absolute difference of the two integrals equals the integral of the (nonnegative) integrand over $A_{2}$ minus the integral over $A_{1}$ in magnitude, and is therefore bounded above by the integral over